• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 01:52:54 on September 19, 2010

    Entry number 332562

    keyword=Calorimeter / DAQ status

    Calorimeter is on with HV at 750V on first crate and 800 V on second crate and 600 V on cosmics paddles. LED is unplugged. Cosmics trigger was setup and will be runnning tonight for the top crate.